OpenStudy (chihiroasleaf):

\[\left(5+6\right)\left(5^2+6^2\right)\left(5^4+6^4\right)\left(5^8+6^8\right)\left(5^{16}+6^{16}\right)\left(5^{32}+6^{32}\right)=\left(6^x-5^y\right)\] \[x-y=...\]

OpenStudy (anonymous):

Multiply the LHS by \(1=(6-5)\), then you have \[\begin{align*}(6-5)\prod_{k=0}^5\left(5^{2k}+6^{2k}\right)&=(6-5)(6+5)\prod_{k=1}^5\left(5^{2k}+6^{2k}\right)\\\\&=(6^2-5^2)\prod_{k=1}^5\left(5^{2k}+6^{2k}\right)\end{align*}\] From here you can factor out the next term in the product: \[\begin{align*}(6^2-5^2)\prod_{k=1}^5\left(5^{2k}+6^{2k}\right)&=(6^2-5^2)(6^2+5^2)\prod_{k=2}^5\left(5^{2k}+6^{2k}\right)\\\\&=(6^4-5^4)\prod_{k=2}^5\left(5^{2k}+6^{2k}\right)\end{align*}\] and so on.

OpenStudy (chihiroasleaf):

thank you... I got x = y = 64 so, x-y =0

OpenStudy (anonymous):

Yes, I checked with Mathematica and got the same result.